OpenStudy (anonymous):

Given that f(x)=1−x^3 is f(f)= 1-(1-x^3)

OpenStudy (turingtest):

No f(x)=1-x^3 f(f)=f[f(x)]=f(1-x^3)=1-(1-x^3)^3 think about that; should makes sense